OpenStudy (anonymous):

The graph of f(x) = x^2 is shown below. Determine the coordinates of point A after the transformation to g(x) = f(x) - 2.

ganeshie8 (ganeshie8):

where is point A

OpenStudy (anonymous):

i need help with this ?

OpenStudy (anonymous):

point A is at 1,1

ganeshie8 (ganeshie8):

g(x) = f(x) - 2 plugin x = 1

ganeshie8 (ganeshie8):

g(x) = f(1) - 2 = 1^2 - 2 = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so x is -1

ganeshie8 (ganeshie8):

yes so the new point wud be (1, -1)

ganeshie8 (ganeshie8):

nop, x is still 1, y is -1.

ganeshie8 (ganeshie8):

hence the new point wud be (1, -1)

OpenStudy (anonymous):

how thats not a answer choice on my assignment

ganeshie8 (ganeshie8):

wat choices do u have

OpenStudy (anonymous):

1, 1) (-1, 1) (3, 1) (2, 1)

ganeshie8 (ganeshie8):

check the given point A once if it is correct

OpenStudy (anonymous):

how

ganeshie8 (ganeshie8):

check if the A is (1, 1) or not in your question

OpenStudy (anonymous):

A is (1,1)

OpenStudy (anonymous):

??

ganeshie8 (ganeshie8):

then (1, -1) is the answer.